OpenStudy (anonymous):

What is the equation of the line passing through (-7,4) whose intercepts are numerically equal but of opposite signs?

OpenStudy (agl202):

Need help?

OpenStudy (anonymous):

Yeah

OpenStudy (agl202):

for this im going to assume you know qudratics y-int = -x x-int = x so we have the points (x,0), (-x,0), (-7,4) so we have gradients of -(4+x)/7 and -4/(7+x) as they are the same gradient -(4+x)/7 = -4/(7+x) solving for -(4+x)/7 = -4/(7+x) (x+4)(x+7)=28 x^2+11x+28=28 x^2+11x=0 x(x+11) x=-11 or x=0 x=0 isn't applicable because of some reason that i dont know so subbing x=-11 in our first cordinates we get (-11,0),(0,11),(-7,4) gradient= (4-11)/(-7-0) = 1 using one point formula y-11=x x-y+11=x or y=x+11

OpenStudy (agl202):

Sorry took a while :( hope it helps!!!:)

OpenStudy (anonymous):

It's alright thank you!

OpenStudy (agl202):

I'm glad I helped dear!!! : )